OpenStudy (anonymous):

Find the standard deviation. Round to one more place than the data. 6, 5, 9, 13, 18, 7, 5, 20, 11, 15 A. 5.4 B. 1.2 C. 4.9 D. 5.1

OpenStudy (anonymous):

first find the mean \[\large x=\frac{6+5+9+13+18+7+5+20+11+15}{10}=10.9\] now for standard deviation \[\Large \sigma=\sqrt{\frac{1}{n}\Sigma _{j=1}^{n}(xj-x)^2}\]

OpenStudy (anonymous):

so using the above values can you find the standard deviation now ?

OpenStudy (anonymous):

standard deviation ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

what is standard deviation

OpenStudy (anonymous):

\[\Large \sigma=5.1\] D

OpenStudy (anonymous):

@85295james

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

did you get it?

OpenStudy (anonymous):

no but i am thinking

OpenStudy (anonymous):

ok let me explain in detail

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

are you okay with the mean ? 10.9 ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

6+ 5+ 9+ 13+18+ 7+ 5+ 20+ 11+ 15 /10=10.9

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

but what do i do with 5.1

OpenStudy (anonymous):

i don't understand

OpenStudy (anonymous):

i am going to find the standard deviation

OpenStudy (anonymous):

did you get it now?

OpenStudy (anonymous):

1.2

OpenStudy (anonymous):

O 5.16

OpenStudy (anonymous):

it is D

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

thank you